# mode=buildnew

### v1

@@ II Collisional Model
We will perform our analysis based on the steady state $\sigma_{S}^{*}$ of the system associated with the map above. To do so, note that the evolution of the in-between system in one time step is always given by
$\sigma_{S}\mapsto\text{GTr}_{A}\left[S_{1}\left(\left(X\sigma_{S}X^{T}+Y\right)\otimes\sigma_{A}^{0}\right)S_{1}^{T}\right],$ (14)
with GTr denoting the Gaussian partial trace. This is a consequence of the fact that the system always interacts with a fresh ancilla from the chain, in the initial state $\sigma_{A}^{0}$. The steady state $\sigma_{S}^{*}$ is obtained as the fixed point of this matrix evolution equation:
$\sigma_{S}^{*}=\text{GTr}_{A}\left[S_{2}\left(\left(X\sigma_{S}^{*}X^{T}+Y\right)\otimes\sigma_{A}^{0}\right)S_{2}^{T}\right]$ (15)
The system is now allowed to evolve with $N$ steps of the collisional model. After this, the state of the system is given by
$\sigma^{N}=\mathcal{U}_{N}\circ\mathcal{E}\circ\dots\circ\mathcal{U}_{1}\circ\mathcal{E}\;\sigma_{0}$ (16)
The final state of the $N$ ancillae, denoted $\Sigma^{N}$, is obtained by performing the partial trace over the degress of freedom of the system:
@@ III.2 Gaussian metrology
$\mathcal{D}(W)\ket{L_{\theta}}=\ket{\partial\sigma_{\theta}}.$ (23)
@@ IV.1 Squeezed ancillae thermometry
For simplicity, we assume that $\varphi=0$ so that $\sigma_{A}^{0}(r,\varphi)=\sigma_{A}^{0}(r):=\text{diag }(e^{2r},e^{-2r})$, since the phase rotations here can be disregarded Mirkhalaf et al. (2023). To further simplify our analysis, we also assume that no TMS is present, so we may take $h=0$ in Eq. (8), dispelling Eq. (5) from the dynamics. Additionally, we keep the coupling strength $\gamma\tau_{SE}$ of the open evolution fixed for all simulations throughout the rest of this section.
@@
For that, we now describe the steps we follow in order to obtain the QFI for this model, summarizing the formalism from Secs. II and  III.3. First, once all the parameters are chosen, we first obtain the steady-state of the system, given by Eq. (15). After all the appropriate maps are applied, we can compute the collective CM of $N$-ancillae, per Eq. (17). This will be the state of interest for the simulations. The reason is that we can now compute the symplectic eigenvalues of the corresponding CM in Eq. (17), leading us to the diagonal matrix $W$ in Eq. (25). Once that is done, it is then trivial to compute the QFI for the $N$-ancilla state by using Eq. (27).
@@ IV.2 Two-mode squeezing thermometry
We start by fixing the BS coupling if Fig. 5 (a), where we plot the QFI density as a function of $N$ for different values of $h\tau_{SA}$. We obtain a very similar behavior to the one observed in Fig. 4, where we can see that the effects are even more pronounced. The solid line shows the numerical fit and the markers correspond to the numerical values. For the $h\tau_{SA}=1$ (red downward markers) we obtain a large value for the fitting parameter, at $\alpha\approx 212$. Note how this is an order of magnitude larger than what we get for most of the other configurations and how this is very clear visually. We can see how the number of ancillae necessary to saturate to its asymptotic value is also much larger, around the order of $10^{3}$, while this happens around $10$ or $10^{2}$ for the other curves. Also note how we can easily perform these simulations even for $N$ or the order of $10^{3}$, showcasing the usefulness of the Gaussian formalism and the compact expression from Eq. (27). Obtaining similar plots for the qubit based model would be out of reach in generic scenarios.
@@ Appendix A Stability of the steady state
Eq. (35) can be solved for its steady state. First, for two any matrices $A$ and $B$, let us define the linear superoperator $A\otimes B$ by $(A\otimes B)C=ACB^{T}$, where $X$ is any other matrix. Then Eq. (35) is, in this language, given by
$\begin{split}\sigma_{S}^{n}&=(X\otimes X)(A\otimes A)\sigma_{S}^{n-1}+(X\otimes X)(B\otimes B)\sigma_{A}+Y\\ \end{split}$ (36)
@@
$G=(X\otimes X)(A\otimes A),\quad H=(X\otimes X)(B\otimes B)\sigma_{A}+Y.$ (38)

### v2

@@ II Collisional Model
We will perform our analysis based on the steady state $\sigma_{S}^{*}$ of the system associated with the map above. To do so, we first introduce the Gaussian partial trace (over a covariance matrix). This operation is defined as:
$\text{GTr}_{A}\;\begin{matrix}\begin{pmatrix}\sigma_{S}&&\sigma_{SA}\\ \sigma_{SA}^{T}&&\sigma_{A}\end{pmatrix}\end{matrix}:=\sigma_{S},$ (14)
given the subsystems A and S and their correlation block $\sigma_{SA}$ Serafini (2017). Now, note that the evolution of the in-between system in one time step is always given by
$\sigma_{S}\mapsto\text{GTr}_{A}\left[S\left(\left(X\sigma_{S}X^{T}+Y\right)\oplus\sigma_{A}^{0}\right)S^{T}\right].$ (15)
This is a consequence of the fact that the system always interacts with a fresh ancilla from the chain, in the initial state $\sigma_{A}^{0}$. The steady state $\sigma_{S}^{*}$ of the system is obtained as the fixed point of this matrix evolution equation:
$\sigma_{S}^{*}=\text{GTr}_{A}\left[S\left(\left(X\sigma_{S}^{*}X^{T}+Y\right)\oplus\sigma_{A}^{0}\right)S^{T}\right]$ (16)
The system is now allowed to evolve with $N$ steps of the collisional model. After this, the collective state of the system + ancilla is given by:
$\sigma^{N}=\mathcal{U}_{N}\circ\mathcal{E}\circ\dots\circ\mathcal{U}_{1}\circ\mathcal{E}\;\sigma_{0}^{*},$ (17)
with
$\sigma_{0}^{*}:=\sigma_{S}^{*}\oplus\sigma_{A}^{0}\oplus\sigma_{A}^{0}\oplus\dots\oplus\sigma_{A}^{0}.$ (18)
Here we have just replaced the (unimportant) initial state of the system with its steady-state $\sigma_{S}^{*}$ from Eq. (17). The final state of the $N$ ancillae, denoted $\Sigma^{N}$, is obtained by performing the partial trace over the degree of freedom of the system:
@@ III.2 Gaussian metrology
$\mathcal{D}(\sigma_{\theta})\ket{L_{\theta}}=\ket{\partial\sigma_{\theta}}.$ (25)
@@ IV.1 Squeezed ancillae thermometry
For simplicity, we assume that $\varphi=0$ so that $\sigma_{A}^{0}(r,\varphi)=\sigma_{A}^{0}(r):=\text{diag }(e^{2r},e^{-2r})$, since the phase rotations here can be disregarded Mirkhalaf et al. (2023). To further simplify our analysis, in this section we also assume that no TMS is present, so we may take $h=0$ in Eq. (8), dispelling Eq. (5) from the dynamics. We analyze the role of the TMS in Sec. IV.2. Additionally, we keep the coupling strength $\gamma\tau_{SE}$ of the open evolution fixed for all simulations throughout the rest of this section.
@@
For that, we now describe the steps we follow in order to obtain the QFI for this model, summarizing the formalism from Secs. II and  III.3. First, once all the parameters are chosen, we first obtain the steady-state of the system, given by Eq. (16). After that we can apply the stroboscopic map (13) to this steady-state, per Eq. (17), obtaining the collective state of the sytem + ancillae. Thus, after all the appropriate maps are applied, we can compute the collective CM of $N$-ancillae by simply tracing out the system, per Eq. (19). This will be the state of interest for the simulations. The reason is that we can now compute the symplectic eigenvalues of the corresponding CM in Eq. (19), leading us to the diagonal matrix $W$ in Eq. (27). Once that is done, it is then trivial to compute the QFI for the $N$-ancilla state by using Eq. (29).
@@
We also make a brief comment regarding the goodness of this fit. One can see that the fit might get slightly worse depending on the choice of parameters, but is unclear to us if or when this fit breaks down. We could not find a combination of parameters where this happens, so this is an open question left to more detailed investigations. More specifically, one could start by exploring different values of $\gamma\tau_{SE}$, which was kept fixed in our simulations. Also, the fit might be a bit unstable numerically whenever $f_{N}$ is approximately constant, since certain numerical routines might be unable to distinguish whether $f_{\infty}\approx f_{1}$ or if $\alpha\rightarrow 0$.
@@ IV.2 Two-mode squeezing thermometry
We start by fixing the BS coupling if Fig. 5 (a), where we plot the QFI density as a function of $N$ for different values of $h\tau_{SA}$. We obtain a very similar behavior to the one observed in Fig. 4, where we can see that the effects are even more pronounced. The solid line shows the numerical fit and the markers correspond to the numerical values. For the $h\tau_{SA}=1$ (red downward markers) we obtain a large value for the fitting parameter, at $\alpha\approx 212$. Note how this is an order of magnitude larger than what we get for most of the other configurations and how this is very clear visually. We can see how the number of ancillae necessary to saturate to its asymptotic value is also much larger, around the order of $10^{3}$, while this happens around $10$ or $10^{2}$ for the other curves. Also note how we can easily perform these simulations even for $N$ or the order of $10^{3}$, showcasing the usefulness of the Gaussian formalism and the compact expression from Eq. (29). Obtaining similar plots for the qubit based model would be out of reach in generic scenarios. For instance, in Ref. Seah et al. (2019) the authors investigate the scaling up to $12$ qubits.
@@ Appendix A Stability of the steady state
Eq. (37) can be solved for its steady state. First, for two any matrices $A$ and $B$, let us define the linear superoperator $A\>\tilde{\otimes}\>B$ by $(A\otimes B)C=ACB^{T}$, where $C$ is any other matrix. Then Eq. (37) is, in this language, given by
$\begin{split}\sigma_{S}^{n}&=(X\>\tilde{\otimes}\>X)(A\>\tilde{\otimes}\>A)\sigma_{S}^{n-1}+(X\>\tilde{\otimes}\>X)(B\>\tilde{\otimes}\>B)\sigma_{A}+Y\\ \end{split}$ (38)
@@
$G=(X\>\tilde{\otimes}\>X)(A\>\tilde{\otimes}\>A),\quad H=(X\>\tilde{\otimes}\>X)(B\>\tilde{\otimes}\>B)\sigma_{A}+Y.$ (40)